# Johan Håstad

### v2

@@ 3.2 Almost complete matchings and restrictions
We let an almost complete matching, usually denoted $\tau$, be an assignment to all variables $y_{P}$ such that exactly half the variables next to any mini-square are true. Note that many such $\tau$ do exist and in particular we can pick half the paths in any group of $2R$ paths to be of the each type. There are many other ways to pick $\tau$ but we do not need this explicitly.
Our restrictions give values to some variables $y_{P}$ and some remain unset. We rename the latter $z_{P}$ to distinguish new and old variables. While an individual $x_{e}$ might depend on many $y_{P}$ it only depends on at most three new variables.
Let us proceed to describe how to pick a random restriction from our space. We make uniformly random choices but if some choice is very unlucky we start all over again. Let $\pi_{1}$ be a fixed matching of all super-squares except that of the designated survivor in the top left corner. We pick a restriction $\sigma$ as follows.
@@
    Pick a uniformly random $\tau$ respecting the condition that half the variables next to any mini-square are true. If any group of $2R$ variables between two fixed mini-squares has fewer than $R/2$ variables of either value, restart. We denote this event as “lopsided group”.
@@
    Pick a random mini-square from each super-square, except the super-square of the designated survivor. Match these mini-squares according to $\pi_{1}$. Hopefully without causing confusion we keep the name $\pi_{1}$ for this matching of mini-squares. These nodes jointly with the designated survivor are the chosen mini-squares and we call this set $U$.
@@
    For each pair of mini-squares $(s_{1},s_{2})$ matched in $\pi_{1}$ pick one augmenting path of type 1 and convert it to type 0. These are called chosen paths. The choice is based on an advice string $B$ as defined below.
@@
    For each pair of mini-squares $(s_{1},s_{2})$ in $U$ in adjacent super-squares but not matched in $\pi_{1}$ pick one augmenting path of type 0 and make it a chosen path. The choice of the path is based on the advice string as defined below. This is also done with $s_{1}$ being the designated survivor.
Let us see that this defines a smaller instance of PHP in a set of new variables $z_{P}$ corresponding to the value of $y_{P}$ on all chosen paths.
@@ 3.3 The reduced instance
By Lemma 3.2 if exactly one variable on a chosen path next to a mini-square in $U$ is true, then it is possible to find a matching of this mini-square. Thus the local conditions of a mini-square turn in to an axiom of the new instance of PHP. Let us see how to replace the old variables in a supposed proof with these new variables. First note that old variables not on edges in chosen mini-squares or in bricks with at least one chosen path are now fixed in a way respecting the corresponding axioms.
Consider a brick with at least one chosen path going through. There is only one chosen path between any two adjacent chosen mini-squares and it is not difficult to see that at most one chosen path goes through any brick. There are two possible matchings of this brick depending on the value of the corresponding variable $z_{P}$. If an edge $e$ is present in neither we replace $x_{e}$ by 0 and if it is present in both we replace it by 1. If $e$ is present in only one we replace it by $z_{P}$ or $\bar{z}_{P}$ in the natural way. It is easy to see that any axiom related to a node in the brick becomes true.
Similarly in a mini-square we have four (or fewer if it is on the perimeter) chosen paths next to it. These are controlled by four new variables that we here locally call $z_{i}$ for $1\leq i\leq 4$. The new local axiom is that exactly one of these four variables is true.
We have four different matchings of the mini-square depending on which $z_{i}$ is chosen to be true. Look at an edge, $e$ and suppose it appears in the matchings corresponding to $z_{2}$ and $z_{3}$ being true. In such a case we replace $x_{e}$ by $z_{2}\lor z_{3}$ and similarly in other cases. If $e$ is in none of the four matchings we replace $x_{e}$ by the constant 0 and if it is in all four we replace it by the constant 1.
It is easy to check that any original axiom inside the mini-square either reduces to true or that exactly one of the four $z_{i}$ is true. Indeed looking at the disjunctions replacing the four variables $x_{e}$ around any node, each $z_{i}$ appears in exactly one. We summarize the discussion of this section as follows.
Lemma 3.3
A restriction $\sigma$ reduces an axiom in the PHP in the grid of size $n\times n$ either to the constant true or an axiom in the PHP of the new variables $z_{P}$. Each variable $x_{e}$ is substituted by a conjunction of up to three literals.
@@
As in previous papers these full restrictions are not used in the main argument and we work with partial restrictions that we now turn to.
@@ 3.4 Partial restrictions
A full restriction is a very rigid object with exactly one live mini-square in each super-square and we proceed to make a more random looking object.
@@
    Pick, without replacement of mini-squares, $k$ uniformly random pairs of mini-squares in adjacent super-squares. These are picked one at the time and each is picked uniformly from the set of remaining possible pairs. This yields a matching $\pi_{2}$. If this pick is unbalanced as defined below redo this step.
@@
    Change the type of one augmenting path between any pairs of nodes in $\pi_{2}$ from 1 to 0. The choice of which of the at least $R/2$ paths of type 1 is based on the advice string and defined below.
The mini-squares picked during this process jointly with the chosen mini-square are considered “alive”. For any path, $P$ between two live mini-squares the corresponding value, $y_{P}$ is now considered undetermined while other $y_{P}$ variables are now fixed. It is not true that these undetermined variables can vary freely as in fact for each pair of live mini-square we may change the value of at most one of its adjacent paths from 0 to 1. No $y_{P}$ will be changed from 1 to 0. In spite of this we do not consider such a value to be known unless we have the full information of one of its end points.
As we pick $2k$ mini-squares we expect roughly [^3] The reason this is not exactly true is super-squares at the perimeter have only two or three neighboring super-squares. Such squares are less likely to have many live mini-squares. This results in a factor $(1+o(1))$ more mini-squares in other super-squares but this small factor does not matter and we ignore it. $2C\log n$ live mini-squares in any super-square. If this number is larger than $4C\log n$ for any super-square then we consider the pick unbalanced and we restart. For any two adjacent super-squares $S_{1}$ and $S_{2}$ we expect $C\log n/2$ pairs $(s_{1},s_{2})$ picked such that $s_{i}\in S_{i}$. If this number is smaller than $C\log n/4$ for any pair $(S_{1},S_{2})$ we consider this unbalanced and restart.
@@ 3.5 Changing types of augmenting paths
In the above procedure, in two places we need to select an augmenting path and (possibly) change its type. This happens when changing a path from type 1 to type 0 because its end-points are matched in $\pi_{1}$ or $\pi_{2}$ and when opening up for changing the type from 0 to 1 by making it a chosen path.
@@
The choices of the $2R$ variables in a group corresponds to a vector in $\{0,1\}^{2R}$ and we want to modify one coordinate in order to change the Hamming weight from $t$ to either $t+1$ or $t-1$ and let us suppose the latter. We want the choice to be limited and as invertible as possible. As the number of strings of weight $t$ and $t-1$ are different we cannot achieve perfect invertability. Suppose first that $t\leq R$.
Definition 3.4
Suppose $t\leq R$. A mapping $f$ mapping (${{2R}\choose t}$ to (${{2R}\choose{t-1}}$ and which maps each set to a subset, is a $k$-almost bijection if it is surjective and $|f^{-1}(x)|\leq k$ for any $x$.
@@
Proof:  Consider a bipartite graph where the the left hand side elements are subsets of size $t-1$ and the right hand side elements are subsets of size $t$. Connect two sets iff one is a subset of the other. It is well known (see for instance Corollary 2.4 in [4]) that this graph has a matching, $M_{1}$, of size ($2R\choose t-1$.
Modify the construction by making three copies of each left hand side node. Each copy is again connected to any set that contains it. It follows by the LYM inequality (stated as Theorem 3.3 in [4]) that this graph has a matching $M_{2}$ of size ($2R\choose t$.
@@
Taking the complement of both input and output we define a $4$-almost bijection, $g$ mapping ($2R\choose t$ to ($2R\choose{t+1}$ for $R\leq t\leq 3R/4$. We use $f$ and $g$ to guide our choices and in addition we have two bits of advice for each group.
Definition 3.6
@@
To make the situation uniform we have two advice bits for any pair of mini-squares in adjacent super-squares. Thus most of these bits are never used. We let $B$ denote the values of all these bits.
If $t\leq R$ we instead consider $f(v)$ and change the type of the corresponding path. Finally if we want to allow to change the weight from $t$ to $t+1$ we reverse the two cases.
@@ 3.6 Analyzing the probability of a restart
Proof:  The almost complete matching $\tau$ is defined by the variables $y_{P}$ which we in this section choose to take values 1 and $-1$. For a uniformly random assignment to all variables, let $Z$ be the vector of all mini-square sums. Let us denote the number of mini-squares by $d$ making $Z$ an integer vector of length $d$. When constructing $\tau$ we are conditioning on the event $Z=0^{d}$.
Fix any two mini-squares $s_{1}$ and $s_{2}$ and let $g$ be the group of $2R$ variables associated with paths between $s_{1}$ and $s_{2}$. Let $X_{g}$ be sum of the variables in this group. We want to estimate the probability that $X_{g}=x$ where $x$ is either at least $R$ or at most $-R$. Let $Z^{\prime}_{g}$ be the set of mini-square sums when the paths between $s_{1}$ and $s_{2}$ are removed. As these two mini-squares also have other adjacent augmenting paths this is still a vector length $d$. Let $v_{x}$ be the vector of length $d$ that has $x$ at positions $s_{1}$ and $s_{2}$ and is otherwise 0. We want to estimate
@@
Lemma 3.8
For any outcome $v\in\mathbb{Z}^{d}$ we have $Pr[Z_{g}^{\prime}=v]\leq Pr[Z_{g}^{\prime}=0^{d}]$. The similar statement is true for any multi-graph where each edge appear an even number of times.
@@
for the probability we want to estimate. Clearly $Pr[Z=0^{d}]\geq Pr[X_{g}=0]Pr[Z_{g}^{\prime}=0^{d}]$ and substituting this into the equation we get the upper bound $Pr[X_{g}=x]/Pr[X_{g}=0]$. When $|x|\geq R$ this probability is $2^{-cR}$ for some explicit $c$ and since there are at most $n^{2}$ pairs of mini-squares the lemma follows.
@@
Proof:  Let $f(v)$ be the probability that $Z^{\prime}_{g}=v$. As $v$ is the vector sum of contributions of single $y_{P}$, $f(v)$ is a giant convolution. A typical term, $f_{0}$, is a probability distribution that give weight $\frac{1}{2}$ to the vector $v_{1}$ and weight $\frac{1}{2}$ to the vector $-v_{1}$ where $v_{1}$ is similar to the vector $v_{x}$ in the above proof. Consider the corresponding Fourier expansion
$\hat{f}_{0}(x)=\sum_{v}f_{0}(v)e^{2\pi i(v,x)}$
where $(v,x)$ is the inner product of $v$ and $x$ and $x$ belongs to the $d$-dimensional torus. As $f_{0}$ is symmetric under negation, $\hat{f}_{0}$ is real-valued. Summing all variables in a given group gives a distribution function which is a $2R$-fold convolution of $f_{0}$. Its Fourier expansion represents $\hat{f}_{0}^{2R}$ which is a positive function. Summing the contribution from all edges corresponds, on the Fourier side, of taking the product. We conclude that
$\hat{f}(x)=\sum_{v}f(v)e^{2\pi i(v,x)}$
is a real-valued positive function. The largest Fourier coefficients of such a function is the constant coefficient and this is exactly what we wanted to prove.
@@
Proof:  As this is a very standard argument let us only sketch it. If we picked the pairs of mini-squares with replacement the lemma would be completely standard. Let us analyze the dynamic process. To see that we do not pick more than $4C\log n$ mini-squares in any super-square with high probability we note two facts.
[…]
@@
We turn to the condition that we have at least $C\log n/4$ pairs in any two adjacent super-squares. Also this analysis is completely standard if edges are picked with replacement. If we condition on not picking more than $4C\log n$ mini-squares in any super-square the probability that a picked edge is between two given super-squares does not decrease by more than a factor $1-o(1)$. Hence the probability of picking very few edges between two given super-squares in the process without replacement is not so different compared to the probability of the same event in the process with replacement. We leave it to the reader to fill in the details.

### v3

@@ 3.2 Almost complete matchings and restrictions
We let an almost complete matching, usually denoted $\tau$, be an assignment to all variables $y_{P}$ such that exactly half the variables next to any mini-square are true. Note that many such $\tau$ do exist and in particular we can pick half the paths in any group of $2R$ paths to be of each type. There are many other ways to pick $\tau$ but we do not need this explicitly.
Our restrictions are constructed by giving fixed values to some variables $y_{P}$ while some remain unset. We rename the latter $z_{P}$ to distinguish new and old variables. While an individual $x_{e}$ might depend on many $y_{P}$ it only depends on at most three new variables.
Let us proceed to describe how to pick a random restriction from our space. We make uniformly random choices but if some choice is very unlucky we redo this choice. Let $\pi_{1}$ be a fixed matching of all super-squares except the top left square where the designated survivor is located. We pick a restriction $\sigma$ as follows.
@@
    Pick a uniformly random almost complete matching $\tau$. If any group of $2R$ variables between two fixed mini-squares has fewer than $R/2$ variables of either value, restart. We denote this event as “lopsided group”.
@@
    Pick a random mini-square from each super-square, except the super-square of the designated survivor forming, jointly with the designated survivor, the set $U$. Match these mini-squares according to $\pi_{1}$. Call the resulting matching (now of mini-squares) $\pi_{1}^{U}$.
@@
    For each pair of mini-squares $(s_{1},s_{2})$ matched in $\pi_{1}^{U}$ pick one augmenting path of type $1$ and convert it to type 0. These are called chosen paths. The choice is based on an advice string $B$ as discussed in Section 3.5 below.
@@
    For each pair of mini-squares $(s_{1},s_{2})$ in $U$ in adjacent super-squares but not matched in $\pi_{1}^{U}$ pick one augmenting path of type 0 and make it a chosen path. The choice of the path is based on the advice string as discussed in Section 3.5 below. This is also done with $s_{1}$ being the designated survivor.
For each chosen path $P$ we have corresponding variable $z_{P}$. These give the variables in the reduced instance.
@@ 3.3 The reduced instance
By Lemma 3.3 if exactly one variable corresponding to a chosen path next to a mini-square in $U$ is true, then it is possible to find a matching of this mini-square. Thus the local conditions of a mini-square turn in to an axiom of the new instance of $\operatorname{PHP}$. Let us see how to replace the old variables in a supposed proof with these new variables. First note that old variables not on edges in chosen mini-squares or in bricks with at least one chosen path are now fixed in a way respecting the corresponding axioms.
We now describe how to restrict the variables of $\operatorname{PHP}_{n}$ according to a full restriction $\sigma$. Consider a brick with at least one chosen path going through. There is only one chosen path between any two adjacent chosen mini-squares and it is not difficult to see that at most one chosen path goes through any brick. There are two possible matchings of this brick depending on the value of the corresponding variable $z_{P}$. If an edge $e$ is present in neither we replace $x_{e}$ by 0 and if it is present in both we replace it by $1$. If $e$ is present in only one we replace it by $z_{P}$ or $\bar{z}_{P}$ in the natural way. It is easy to see that any axiom related to a node in the brick becomes true.
Similarly, in a mini-square we have four (or fewer if it is on the perimeter) chosen paths next to it. These are controlled by four new variables that we here locally call $z_{i}$ for $1\leq i\leq 4$. The new local axiom is that exactly one of these four variables is true.
We have four different matchings of the mini-square depending on which $z_{i}$ is chosen to be true. Look at an edge, $e$ and suppose, for example, that it appears in the matchings corresponding to $z_{2}$ and $z_{3}$ being true. In such a case we replace $x_{e}$ by $z_{2}\lor z_{3}$ and similarly in other cases. If $e$ is in none of the four matchings we replace $x_{e}$ by the constant 0 and if it is in all four we replace it by the constant $1$.
It is easy to check that any original axiom inside the mini-square either reduces to true or that exactly one of the four $z_{i}$ is true. Indeed, looking at the disjunctions replacing the four variables $x_{e}$ around any node, each $z_{i}$ appears in exactly one. We summarize the discussion of this section as follows.
Lemma 3.4.
A restriction $\sigma$ reduces an axiom in the $\operatorname{PHP}_{n}$ either to the constant true or an axiom in the $\operatorname{PHP}_{(n/T)}$ of the new variables $z_{P}$. Each variable $x_{e}$ is substituted by a conjunction of up to three literals.
@@
As in previous papers these full restrictions are not used in the main argument and we work with partial restrictions that we now turn to. The intuitive reason to introduce these is that a full restriction is a very rigid object with exactly one live mini-square in each super-square. In the final argument we compare the number of restrictions to the number of restrictions with slightly less live variables. This does not work for full restrictions as the number of full restrictions that have $s$ live mini-squares removed is actually higher than the number of full restrictions. If we instead first globally pick a larger number of mini-squares to keep alive, then decreasing the number of such live mini-squares does decrease the number of alternatives. Let us turn to defining these partial restrictions.
@@ 3.4 Partial restrictions
    Pick, without replacement of mini-squares, $k$ uniformly random pairs of mini-squares in adjacent super-squares. These are picked one at the time and each is picked uniformly from the set of remaining possible pairs. This yields a matching $\pi_{2}$. If any super-square has more than $4C\log n$ live mini-squares we consider the choice unbalanced and we restart. Also, if for any two adjacent super-squares $S_{1}$ and $S_{2}$ if we have less $C\log n/4$ pairs $(s_{1},s_{2})$ with $s_{i}\in S_{i}$ we consider the choice unbalanced and restart.
@@
    Change the type of one augmenting path between any pairs of nodes in $\pi_{2}$ from $1$ to 0. The choice of which of the at least $R/2$ paths of type $1$ is based on the advice string as discussed in Section 3.5 below.
The mini-squares picked during this process jointly with the chosen mini-square are considered “alive”. For any path, $P$ between two live mini-squares the corresponding value, $y_{P}$, is now considered undetermined while other $y_{P}$ variables are fixed. When we later go from a partial restriction to a full restriction it is not true that these undetermined variables can vary freely as in fact for each pair of live mini-square we may change the value of at most one of its adjacent paths. In spite of this we do not consider such a value to be known unless we have the full information at one of its end points.
As we pick $2k$ mini-squares we expect roughly [^3] The reason this is not exactly true is super-squares at the perimeter have only two or three neighboring super-squares. Such squares are less likely to have many live mini-squares. This results in a factor $(1+o(1))$ more mini-squares in other super-squares but this small factor does not matter and we ignore it. $2C\log n$ live mini-squares in any super-square and hence, by standard Chernoff bounds it is unlikely that this number is larger than $4C\log n$ for any super-square. Similarly, for any two adjacent super-squares $S_{1}$ and $S_{2}$ we expect $C\log n/2$ pairs $(s_{1},s_{2})$ picked such that $s_{i}\in S_{i}$ and hence it is unlikely that this number is smaller than $C\log n/4$ for any pair $(S_{1},S_{2})$.
@@ 3.5 Changing types of augmenting paths
In the above procedure, in two places we need to select an augmenting path and (possibly) change its type. This happens when changing a path from type $1$ to type 0 because its end-points are matched in $\pi_{1}^{U}$ or $\pi_{2}$ and when opening up for changing the type from 0 to $1$ by making it a chosen path.
@@
The choices of the $2R$ variables in a group corresponds to a vector in $\{0,1\}^{2R}$ and we want to modify one coordinate in order to change the Hamming weight from $t$ to either $t+1$ or t1t-1 and let us suppose the latter. We want the choice to be limited and as invertible as possible. As the number of strings of weights $t$ and t1t-1 are different we cannot achieve perfect invertability. Suppose first that $t\leq R$.
Definition 3.5.
Suppose $t\leq R$. A mapping $f$ mapping ${2R}\choose t$} to ${2R}\choose{t-1}$} which maps each set to a subset, is a $k$-almost bijection if it is surjective and $|f^{-1}(x)|\leq k$ for any $x$.
@@
Proof 3.7.
Consider a bipartite graph where the left hand side elements are subsets of size t1t-1 and the right hand side elements are subsets of size $t$. Connect two sets iff one is a subset of the other. It is well known (see for instance Corollary 2.4 in [4]) that this graph has a matching, $M_{1}$, of size $2R\choose t-1$.
Modify the construction by making three copies of each left hand side node. Each copy is again connected to any set that contains it. It follows by the LYM inequality (stated as Theorem 3.3 in [4]) that this graph has a matching $M_{2}$ of size $2R\choose t$.
@@
Taking the complement of both input and output we define a $4$-almost bijection, $g$ mapping $2R\choose t$ to $2R\choose{t+1}$ for $R\leq t\leq 3R/2$. We use $f$ and $g$ to guide our choices and in addition we have two bits of advice for each group.
Definition 3.8.
@@
To make the situation uniform we have two advice bits for any pair of mini-squares in adjacent super-squares. One can note that most of these bits are never used but they make the construction uniform. We let $B$ denote the values of all these bits.
If $t\leq R$ we instead consider $f(v)$ and change the type of the corresponding path. Finally if we want to change the weight from $t$ to $t+1$ we reverse the two cases.
@@ 3.6 Analyzing the probability of a restart
Proof 3.10.
The almost complete matching $\tau$ is defined by the variables $y_{P}$ which we in this section choose to take values $1$ and 1-1. For each mini-square, sum the variables on its boundary and for a uniformly random assignment to all variables, let $Z$ be the vector of all these mini-square sums. Let us denote the number of mini-squares by $d$ making $Z$ an integer vector of length $d$. When constructing $\tau$ we are conditioning on the event $Z=0^{d}$.
Fix any two mini-squares $s_{1}$ and $s_{2}$ and let $g$ be the group of $2R$ variables associated with paths between $s_{1}$ and $s_{2}$. Let $X_{g}$ be sum of the variables in this group. We want to estimate the probability that $X_{g}=x$ where $x$ is either at most R-R or at least $R$. Let $Z^{\prime}_{g}$ be the set of mini-square sums when the paths between $s_{1}$ and $s_{2}$ are removed. As these two mini-squares also have other adjacent augmenting paths this is still a vector of length $d$. Let $v_{x}$ be the vector of length $d$ that has $x$ at positions $s_{1}$ and $s_{2}$ and is otherwise 0. We want to estimate
@@
Lemma 3.11.
For any outcome $v\in\mathbb{Z}^{d}$ we have $Pr[Z_{g}^{\prime}=v]\leq Pr[Z_{g}^{\prime}=0^{d}]$.
@@
for the probability we want to estimate. Clearly $Pr[Z=0^{d}]\geq Pr[X_{g}=0]Pr[Z_{g}^{\prime}=0^{d}]$ and substituting this into the equation we get the upper bound $Pr[X_{g}=x]/Pr[X_{g}=0]$. When $|x|\geq R$, then by standard Chernoff bounds, this probability is $2^{-cR}$ for some explicit $c$ and since there are at most $n^{2}$ pairs of mini-squares the lemma follows.
@@
Proof 3.12 (Proof of Lemma 3.11).
Let $f(v)$ be the probability that $Z^{\prime}_{g}=v$. As $v$ is the vector sum of contributions of single $y_{P}$, $f(v)$ is a giant convolution. To be more precise for each path $P$ between mini-squares $s_{1}$ and $s_{2}$ we have a vector $v_{1}$ with a one in positions corresponding to $s_{1}$ and $s_{2}$ and 0 in all other positions. Define a probability distribution $f_{P}$ that gives probability $\frac{1}{2}$ to each of $v_{1}$ and $v_{1}$. The function $f(v)$ is the convolution of $f_{P}$ over all paths $P$. If $\hat{f}_{P}$ is the Fourier transform of $f_{P}$ then the Fourier transform of $f$ is
$\hat{f}(x)=\prod_{P}\hat{f}_{P}(x),$ (1)
where $x$ belong to the $d$-dimensional torus. Note that as $f_{P}$ is symmetric around $0^{d}$, $\hat{f}_{P}$ is real-valued. Moreover, for each pair of neighboring mini-squares $s_{1}$ and $s_{2}$ we have $2R$ paths between $s_{1}$ and $s_{2}$. This implies that the right-hand side of (1) can be written as
$\prod_{s_{1},s_{2}}\hat{f}_{s_{1},s_{2}}^{2R},$
where $f_{s_{1},s_{2}}=f_{P}$ for any path $P$ with endpoints $s_{1}$ and $s_{2}$. We conclude that $\hat{f}$ only takes real and non-negative values. We conclude that, as for any function with a real-valued and non-negative Fourier transform, $f(0)\geq f(v)$ for any $v$ and this is exactly what we wanted to prove.
@@
Proof 3.14.
This might follow from the fact that Chernoff bounds are true for negatively correlated variables. It is, however, not quite obvious that our variables are negatively correlated. It is true that fixing the size of the matching causes negative correlation but the requirement to be a matching is more complicated and could give positive correlations. In view of this let us sketch a direct argument.
If we picked the pairs of mini-squares with replacement the lemma would be completely standard. Let us analyze the dynamic process. To see that we do not pick more than $4C\log n$ mini-squares in any super-square with high probability we note two facts.
  • As we only pick an $o(1)$ fraction of all mini-squares, at each point in time a fraction (1o(1))(1-o(1)) of all pairs are available.
@@
We turn to the condition that we have at least $C\log n/4$ pairs in any two adjacent super-squares. Also, this analysis is completely standard if edges are picked with replacement. If we condition on not picking more than $4C\log n$ mini-squares in any super-square the probability that a picked edge is between two given super-squares does not decrease by more than a factor 1o(1)1-o(1). Hence the probability of picking very few edges between two given super-squares in the process without replacement is not so different compared to the probability of the same event in the process with replacement. We leave it to the reader to fill in the details.